OpenStudy (anonymous):

How to find point of intersection of line y = mx+c and parabola y^2 = 4ax? Any help would be appreciated.

OpenStudy (anonymous):

(mx+c)^2 = 4ax

OpenStudy (anonymous):

substituting y from the line eqn into the parabola

OpenStudy (anonymous):

you mean to say, square the term ( mx+c) and equate it to 4ax?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

youll get a quadratic eqn in x..you solve it and obtain 2 values of x

OpenStudy (anonymous):

then u plug in those values of x into any one of the two eqns and find the 2 corresponding values of y

OpenStudy (anonymous):

ok. I will try it. I never thought it would be so simple. And what abt x^2 = 4ay.

OpenStudy (anonymous):

then plug value of y into the parabola x^2= 4a(mx+c)

OpenStudy (anonymous):

Ok. let me note it down.Thanks for ur help.I appreciate it.

OpenStudy (anonymous):

for finding any point of intersection u just just consider the 2 eqns as 2 variable eqns..and solve for x and y by eliminating one of them...whichever is feasible

OpenStudy (anonymous):

where r ufrom? wht class?

OpenStudy (anonymous):

I am from India.This is BCA maths preparation.

OpenStudy (anonymous):

hey i know ur indian....BCA maths prep ..so ur 18? frgiv my ignorance..i dunno abt bca

OpenStudy (anonymous):

Bachelor in Computer Applications.

OpenStudy (anonymous):

arre yaar that i know/......

OpenStudy (anonymous):

U also seem Indian...

OpenStudy (anonymous):

yes...i just dunno at wht level do u start bca prep..in class 12 or after?

OpenStudy (anonymous):

No, this is correspondence course.U can do it any time.

OpenStudy (anonymous):

ohk....right miss...best of luck....hope we c more of each other here...

OpenStudy (anonymous):

yeah, thanks for ur help.

OpenStudy (anonymous):

no worries helping sweet women: )